Jane invested $50,000 into two accounts. The first paid 7% interest and the second paid 12% interest. If she earned $4350 the first year, how much money did she invest in each account?
Answer by robertb(5830)

0.07x + 0.12(50,000 - x) = 4350,
0.07x + 6000-0.12x = 4350,
-0.05x = -1650,
x = 33,000, the amount of investment at 7%,
50,000 - x = 17,000, the amount of investment at 12%.